### v1

Abstract
Large Language Models (LLMs) exhibit remarkably powerful capabilities. One of the crucial factors to achieve success is aligning the LLM’s output with human preferences. This alignment process often requires only a small amount of data to efficiently enhance the LLM’s performance. While effective, research in this area spans multiple domains, and the methods involved are relatively complex to understand. The relationships between different methods have been under-explored, limiting the development of the preference alignment. In light of this, we break down the existing popular alignment strategies into different components and provide a unified framework to study the current alignment strategies, thereby establishing connections among them. In this survey, we decompose all the strategies in preference learning into four components: model, data, feedback, and algorithm. This unified view offers an in-depth understanding of existing alignment algorithms and also opens up possibilities to synergize the strengths of different strategies. Furthermore, we present detailed working examples of prevalent existing algorithms to facilitate a comprehensive understanding for the readers. Finally, based on our unified perspective, we explore the challenges and future research directions for aligning large language models with human preferences.
1 Introduction
Figure 1: A unified view and an illustrative example of preference learning for LLMs.
Represented by ChatGPT [^1] https://chatgpt.com, the rise of large language models (LLMs) has showcased impressive language capabilities and professional competence, as well as providing correct, polite, and knowledgeable responses, which is surprising and admirable. This performance should be largely attributed to the preference alignment process, a currently necessary step that LLMs undergo before public deployment, to prevent LLMs from potentially generating offensive, toxic, or misleading content.
Although large language models (LLMs) have demonstrated impressive capabilities across various fields (geminiteam2024gemini15unlockingmultimodal; deepseekai2024deepseekllmscalingopensource; yang2024qwen2technicalreport; openai2024gpt4technicalreport), they still face challenges in ethics (jiao2024navigatingllmethicsadvancements), safety (shavit2023practices; kumar2023certifying; wei2024jailbroken), and reasoning (yu2023metamath; lightman2023lets; wang2024mathshepherd). In response, numerous alignment-related initiatives have emerged to better address these issues (ouyang2022training; rafailov2024direct; ethayarajh2024kto; meng2024simpo). The snowballing interest also inspires this survey. While many works (shen2023large; wang2024essence) have extensively discussed the concept of alignment, the relationships among the various algorithms of preference learning remain fragmented, lacking a cohesive framework to unify them. To bridge this gap, we aim to provide a systematic framework for preference alignment, as shown in figure 1. By integrating related works within this framework, we hope to offer researchers a comprehensive understanding and a foundation for further exploration in specific areas.
Traditional categorization perspectives (shen2023large; wang2024essence; jiang2024surveyhumanpreferencelearning) tend to split existing methods into reinforcement learning (RL) based methods, like RLHF (ouyang2022training) which requires a reward model for online RL, and supervised finetuning (SFT) based methods like Direct Preference Optimization (DPO) (rafailov2024direct) that directly employs preference optimization within an offline setting. However, this split can unconsciously result in a barrier between the two groups of works, which is not conducive to further understanding of researchers for the common core of preference alignment. Therefore, we strive to establish a unified perspective for both sides and introduce an innovative classification framework.
{forest}
forked edges, for tree= grow=east, reversed=true, anchor=base west, parent anchor=east, child anchor=west, base=center, font=, rectangle, draw=hidden-draw, rounded corners, align=left, text centered, minimum width=4em, edge+=darkgray, line width=1pt, s sep=3pt, inner xsep=2pt, inner ysep=3pt, line width=0.8pt, ver/.style=rotate=90, child anchor=north, parent anchor=south, anchor=center, , where level=1text width=9em,font=,, where level=2text width=8em,font=, where level=3text width=10em,font=,, where level=4text width=6em,font=,, [ Preference Learning, ver [ Preference Data (§4) [ On-policy [ Top-K/Nucleus Sampling holtzman2020curious, Beam Search graves2012sequence, MCTS KocsisMCTS , leaf, text width=33.6em ] ] [ Off-policy [ Data from Human [ OpenAI’s Human Preference ouyang2022training, HH-RLHF bai2022training,
SHP pmlr-v162-ethayarajh22a, Webgpt nakano2021webgpt , leaf, text width=22.0em ] ] [ Data from LLM [ RLAIF lee2023rlaif, Open-Hermes-Preferences open_hermes_preferences,
ULTRAFEEDBACK cui2023ultrafeedback, UltraChat ding2023enhancing , leaf, text width=22.0em ] ] ] ] [ Feedback (§5), [ Direct Reward [ Math: RFT rft, DeepSeekProver (xin2024deepseekproverv15harnessingproofassistant) (xin2024deepseekproveradvancingtheoremproving)
Translation: CPO xu2024contrastive Summarization: PRELUDE gao2024aligning
Code: Pangu-coder2 shen2023pangucoder2, StepCoder dou2024stepcoder, Rltf liu2023rltf , leaf, text width=33.6em ] ] [ Model-based
Reward [ Reward Modeling [ RLAIF lee2023rlaif, RBoN jinnai2024regularized, West-of-N pace2024westofn,
RM-ensemble coste2024reward, LoRA-ensemble zhai2023uncertaintypenalized,
WARM ramé2024warm, Efficient-ensemble zhang2024improving,
Fine-Grained RLHF NEURIPS2023_b8c90b65, PRM uesato2022solving,  lightman2023lets,
OVM yu2024ovm, MATH-Shepherd wang2024mathshepherd,
Prior contraints RM zhou2024prior, Math-Minos gao2024llmcriticshelpcatch , leaf, text width=22.0em ] ] [ Pair-wise Scoring [ Llm-blender (jiang2023llmblender), PandaLM (wang2023pandalm) , leaf, text width=22.0em ] ] [ LLM-as-a-Judge [ Self-Reward (yuan2024self), Meta-Reward (wu2024meta),
CriticGPT (mcaleese2024llmcriticshelpcatch), Generative Verifier (zhang2024generativeverifiersrewardmodeling) , leaf, text width=22.0em ] ] ] ] [ Optimization (§6) [ Pointwise method [ RFT (rft; star; raft), PPO (schulman2017proximal), GRPO (shao2024deepseekmath), ReMax (li2023remax), KTO (ethayarajh2024kto) , leaf, text width=33.6em ] ] [ Pairwise contrast [ CoH (liu2024chain), SLiC (zhao2022calibrating), DPO (rafailov2024direct), IPO (azar2023general), Sr-DPO (yu2024direct), ORPO (hong2024orpo),
Mallow-DPO (chen2024mallows), GRPO (ramesh2024group), DPO-positive (pal2024smaug), CPL (hejna2023contrastive), EXO (ji2024towards),
SimPO (meng2024simpo), sDPO (kim2024sdpo), TR-DPO (gorbatovski2024learn), RSO (liu2023statistical), $f$-DPO (wang2023beyond), CPO (guo2024controllable),
MAPO (she2024mapo), KnowTuning (lyu2024knowtuning), TS-align (zhang2024ts), MODPO (zhou2023beyond), HPO (badrinath2024hybrid) , leaf, text width=33.6em ] ] [ Listwise contrast [ RRHF (yuan2024rrhf), PRO (song2023preference), CycleAlign (hong2023cyclealign), AFT (wang2023making),
VCB (mao2024don), LiPO (liu2024lipo), LIRE (zhu2024lire) , leaf, text width=33.6em ] ] [ Training-Free [ Input Optimization [ BPO (cheng2023black), URIAL (lin2023unlocking), OPO (xu2023align) , leaf, text width=22.0em ] ] [ Output Optimization [ ICDPO (song2024icdpo), Aligner (ji2024aligner), RAIN (li2023rain),
DecodingControl (yang2021fudge; mudgal2023controlled; deng2023reward; liu2024decoding), DeAL (huang2024deal) , leaf, text width=22.0em ] ] ] ] [ Evaluation (§7) [ Rule-based [ Factuality hendrycks2020measuring, Math cobbe2021training; yu2023metamath, Reasoning suzgun2022challenging,
Closed-Book QA joshi2017triviaqa; kwiatkowski2019natural, Coding austin2021program; chen2021evaluating , leaf, text width=33.6em ] ] [ LLM-based [ G-Eval liu2303g, AuPEL wang2023automated, ICE jain2023multi, GEMBA kocmi2023large,
FairEval wang2023large, Auto-J li2023generative, MT-Bench zheng2024judging, Prometheus kim2023prometheus,
Pandalm wang2023pandalm, PRD li2023prd, LLMBar zeng2023evaluating, LLMEval2 zhang2023wider , leaf, text width=33.6em ] ] ] ]
Figure 2: Taxonomy of Preference Learning.
This new framework pivots on two key insights: First, the distinction between on-policy and off-policy settings essentially depends on different sources of data, which can be decoupled from algorithms like PPO or DPO. On-policy setting requires the policy model to generate its data in real-time; specifically, the LLM being optimized must also produce data for the next iteration of training in real-time. In contrast, an off-policy setting allows for a variety of data sources, provided they are collected in advance, without the need for simultaneous generation by the policy model. Many current works employ the transition of specific algorithms between on-policy and off-policy settings (guo2024direct; shao2024deepseekmath). Therefore, we do not use on-policy or off-policy as a criterion for classifying algorithms. Second, inspired by existing work (shao2024deepseekmath), the essence of the objective of optimization in reinforcement learning and supervised finetuning-based methods are actually quite similar. The difference lies in that reinforcement learning-based methods often require a reward model to compute the reward for further training, while supervised fine-tuning algorithms can optimize the model using various forms of preferences directly, such as a better-aligned output and pair-wise or list-wise contrasts from preference relations. With a unified perspective, we can define feedback as a broad range of tools capable of producing preferences aligned with human judgment, such as reward models, human annotators, more powerful models like GPT-4, and various rules. Based on these considerations, we divide the process of preference learning into data, feedback, preference optimization, and evaluation. The taxonomy of our paper is shown in Figure 2.
In summary, our paper investigates and organizes existing work related to preference learning for LLM, offering a unified and novel perspective. Further, based on the content of this review, we summarize several future directions in this area, intending to bring insights for further research.
Figure 3: Examples of the preference learning. Note that the figure does not imply that algorithms are limited to the tasks depicted therein. Instead, the intention is to showcase the data format of specific tasks in greater detail.
Figure 4: Examples of the preference learning strategies with point-wise loss. Similar to the Figure 3, different methods can be adapted to different tasks.
2 Definition and Formulation
Figure 5: The overview of the preference learning. For an LLM $\pi_{\theta}$ to be aligned with human preferences, first we need to prepare preference data. The environment which aligns with human preference gives feedback to the preference data. Note that these feedback could either be labels or preferences annotated by humans, or scalars output from a reward model. By feeding the model, data, and feedback to a specific algorithm, we obtain a LLM $\pi_{\theta}^{\prime}$ that is aligned with human preferences.
In this section, we begin by providing our definition of preference learning for LLM: Given a distribution of the general human preference $\mathcal{P}(x,y)$, where $x$ is a prompt, $y$ is the corresponding output of the LLM, preference learning for LLM $\pi_{\theta}$ is a paradigm that produces a new LLM $\pi_{\theta}^{\prime}$ that align to $\mathcal{P}(x,y)$, where $\mathcal{P}(x,y_{\theta}^{\prime}(x))>\mathcal{P}(x,y_{\theta}(x))$.
To enable LLMs to learn human preferences, the process often involves providing a data sample with input $x$ and corresponding response $y$, and the environment with human preference $\mathcal{P}(x,y)$ assigning feedback to it. Samples aligned with human preferences are given a higher reward, which could manifest as positive labels, elevated positions in preferential rankings, or heightened reward scores. After obtaining the data, the policy model $\pi_{\theta}^{\prime}$ is optimized by a specific algorithm.
Furthermore, it is necessary to explain the relation between preference learning for LLMs and some related concepts, based on this definition. (1) Alignment: Following kenton2021alignment, alignment refers to the research focuses on tackling the so-called behavior alignment problem: How do we create an agent that behaves in accordance with what a human wants? Based on this definition, we regard preference learning for LLMs as a category of methods aimed at achieving alignment. The scope of this paper is confined to textual preference alignment which doesn’t contain other well-known alignment topics such as hallucination, multi-modal alignment, and instruction tuning. (2) Reinforcement Learning from Human Feedback (RLHF): Different from RLHF, the scope of this paper not only includes RL-based methods but also encompasses the traditional called SFT-based methods. What’s more, we adopt a unified perspective to investigate both reinforcement-learning and supervised-learning-based methods.
3 The Unified View of Preference Learning for LLM
Inspired by recent works shao2024deepseekmath; guo2024direct, we survey existing works from a unified perspective in the following two folds:
First, the optimization objectives of RL and SFT-based methods can be described within the same framework. Following shao2024deepseekmath, the gradient with respect to the parameter $\theta$ of a training method can be written as:
$\nabla_{\theta}=\mathbb{E}_{[(q,o)\sim\mathcal{D}]}\left(\frac{1}{|o|}\sum\limits_{t=1}^{|o|}\delta_{\mathcal{A}(r,q,o,t)}\nabla_{\theta}\log\pi_{\theta}(o_{t}|q,o_{<t})\right),$ (1)
where $\mathcal{D}$ denotes the data source which contains the input question $q$ and the output $o$. $\delta$ denotes the gradient coefficient which directly determines the direction and step size of the preference optimization. $\mathcal{A}$ denotes the algorithm. The gradient coefficient is determined by the specific algorithm, data, and corresponding feedback. Tracing back to one significant source influencing the gradient coefficient is the feedback. Note that the feedback could take on various forms. For example, the correctness of data in RFT rft or the preference label in DPO rafailov2024direct could impact the gradient coefficient, thereby affecting the final gradient. Consequently, we define the feedback in our paper as the preference given by the environment that can affect the gradient coefficient. Notably, both RL-based and SFT-based methods can be encapsulated within this framework.
Second, the algorithm can be decoupled from online/offline settings. In the context of alignment, online learning refers to the preference oracle $r$ or its approximator $\hat{r}$ can be queried over training, i.e. the feedback of the responses sampled from the current actor model can be given on the fly. If the feedback signal cannot be obtained in real-time, then it is considered offline learning.
From a traditional perspective, RL-based methods are more flexible with online/offline settings, while SFT-based methods are typically offline. However, as in the first point where we have unified RL-based and SFT-based approaches, it can be inferred that SFT-based methods can also be applied in an online setting, which has been proved by recent work guo2024direct. Actually, what determines whether the setting is online or offline is merely whether the preference signal is generated in real-time or pre-stored. In section 4, we elucidate the methods of acquiring data for both online and offline settings. In online settings, data collection typically follows an on-policy strategy, while in offline settings, it generally adheres to an off-policy strategy. Although it is feasible to combine online feedback collection with an off-policy strategy, such instances are relatively rare in the existing works. Therefore, unlike the categorization in other survey papers wang2023aligning; shen2023large; wang2024essence, we do not use online/offline nor RL/SFT as criteria for classifying algorithms. Instead, we decouple the algorithm from the online/offline setting. As an example, the DPO algorithm does not necessarily have to be offline. It depends on the context in which they are actually applied. If there is an evaluator available to assess preference relations in real-time generated data, then DPO can also be utilized for online optimization.
Based on the two points discussed above, we ultimately divide preference learning into four key elements: Model, Data, Feedback, and Algorithm, as shown in Figure 5.
The process of preference learning can be described as follows: For a LLM $\pi_{\theta}$ to be aligned, we first need to prepare the data $\mathcal{D}$ for training. If we are in an online setting, we must sample behavioral data from the model and the environment will provide a preference feedback signal to the data in real-time. If it is an offline setting, we need to have a preference dataset prepared in advance. Whether it conforms to human preference will be reflected in the feedback $\mathcal{R}$. For example, in the DPO series of methods, data that do not meet the preferences will be assigned a bad label. In RFT, it will be discarded, that is, the gradient coefficient will be zero. For RL-based algorithms like PPO, this would correspond to a lower reward score. Subsequently, the tuple $(\mathcal{D}_{\{x,y\}},\mathcal{R},\pi_{\theta})$ is fed into the algorithm $\mathcal{A}$. We categorize algorithms into four types based on the data required for each model update by the algorithm: training-free methods, point-wise methods, pair-wise contrasts, and list-wise contrasts, without the need to be concerned whether they are RL or SFT-based algorithms. Finally, we acquire an aligned LLM $\pi_{\theta}^{\prime}$. The formal description of this process is provided in Algorithm 1.
1
Input : $\pi_{\theta}$ (Initialize LLM to be aligned), $\mathcal{E}$ (Environment with human preference),
$\mathcal{Q}$ (Unlabeled queries) or $\mathcal{D}$ (Pre-prepared offline dataset), $\mathcal{A}$ (Algorithm)
Output : $\pi_{\theta}^{\prime}$ (Aligned LLM)
2 if reference model is needed then
3       $\pi_{ref}\leftarrow\pi_{\theta}$;
4
5 end if
6
7while (Total training steps not reached) do
8       if online setting then
9             $\mathcal{B}\leftarrow$ Sample response from $\pi_{\theta}$ using $\mathcal{Q}$ and get the preference siganl in real time;
10       end if
11      else if offline setting then
12             $\mathcal{B}\leftarrow$ Get a batch of data with preference siganl from the pre-stored $\mathcal{D}$;
13
14       end if
15      $\pi_{\theta}^{\prime}\leftarrow$ Feed ($(\mathcal{B}_{\{x,y\}},\mathcal{R},\pi_{\theta},\pi_{ref})$ into $\mathcal{A}$ and update model;
16       $\pi_{\theta}\leftarrow\pi_{\theta}^{\prime}$;
17
18 end while
19
20return $\pi_{\theta}$;
21 $\triangleright$ Return the aligned LLM
Algorithm 1 Preference Learning
4 Preference Data
The preference data does not have a fixed form and we use the simplest notation to represent preference data as $(x,y,r)$. Here $x,y$ are the literal information input and the candidate output. $r$ is a preference label given by certain feedback systems, which could be human, reward models, or other scoring systems.
Current LLM preference learning methods gather the training data from two sources: on-policy or off-policy. Generally speaking, the on-policy data collection means we collect the data directly from our policy LLM $\pi_{\theta^{t}}$ at each training step $t$. The off-policy data collection could be done outside the box, independent from the LLM to conduct preference learning and result in a dataset consisting of data that is not generated by the policy model itself. Notably, using preference data sampled from $\pi_{\theta^{0}}$ to train $\pi_{\theta^{t}}$ for $t>0$ is also off-policy.
4.1 On-policy Data Collection
On-policy data collection process is similar to the setting of on-policy reinforcement learning, where the preference data is obtained directly during training: it first samples a batch of experience by the policy LLM and then obtains the reward by interacting with the environment and finally uses it to update the policy LLM. Under such conditions, different methods vary in the preference generator $g(x)$ from the environment.
On-policy Sampling methods
To sample various experiences from the environment, numerous research studies have explored different strategies for decoding.
Diverse sampling strategies such as Top-K/Nucleus Sampling holtzman2020curious and Beam Search graves2012sequence are employed during the generation process of LLMs. These methods determine the efficiency and effectiveness of the data used for preference learning.
For problems that involve multi-step solutions, there has also been some research (wang2024mathshepherd; zhang2024accessinggpt4levelmathematical; zhang2024restmctsllmselftrainingprocess; feng2024alphazeroliketreesearchguidelarge; luo2024improvemathematicalreasoninglanguage; qi2024mutualreasoningmakessmaller; zhu2024recoveringmentalrepresentationslarge) employing Monte Carlo tree search (MCTS) KocsisMCTS to enhance the diversity and performance of the data sampling. MCTS originated from the advancements made in AlphaGo. The fundamental concept of MCTS involves evaluating various strategies through numerous simulations, or rollouts, to determine which strategy yields superior results. This approach resembles a methodical and deliberative thinking process, contrasting with greedy decoding methods that prioritize immediate gains. The core operations of MCTS can be categorized into four distinct phases: selection, expansion, simulation, and backpropagation. MCTS’s efficient search strategy enables models to generate higher-quality data while simultaneously acquiring step-level labels. This refined data can subsequently be utilized to enhance model performance during decoding (qi2024mutualreasoningmakessmaller; zhu2024recoveringmentalrepresentationslarge; feng2024alphazeroliketreesearchguidelarge), train reward models (wang2024mathshepherd; luo2024improvemathematicalreasoninglanguage), and fine-tune the model (zhang2024accessinggpt4levelmathematical; zhang2024restmctsllmselftrainingprocess; feng2024alphazeroliketreesearchguidelarge).
4.2 Off-policy Data Collection
Off-policy data collection entails gathering training data independently of the LLM’s learning process. This method is generally easier than on-policy data collection, largely due to the availability of open-source preference datasets. Alternatively, we can also compile a dataset in advance using the initial model $\pi_{\theta^{0}}$. The off-policy data collection strategy ensures a more diverse training dataset that can often yield improvements in the LLM’s preference learning process. There are two main sources of preference data, the ones from human annotators and those generated by more advanced LLMs.
Data from Human
OpenAI’s Human Preferences ouyang2022training originates from a carefully selected portion of Reddit’s TL;DR corpus volske-etal-2017-tl. Each entry within this dataset consists of a post, paired with two alternative summary options, and is augmented by an evaluation from a human annotator who designates the preferred summary of the two.
HH-RLHF bai2022training involves 170K chats between humans and AI helpers. In these chats, the AI gives two different replies. A human annotator marks which reply is better and which is not as good.
SHP pmlr-v162-ethayarajh22a collects 385K human preferences over responses to questions/instructions in 18 domains for training RLHF reward models and NLG evaluation models.
Webgpt nakano2021webgpt has 20K comparisons, where each example consists of a question, a pair of model answers, and human-rated preference scores for each answer.
Data from LLMs
Obtaining preference from human could be resource-consuming. However, researches (lee2023rlaif; cui2023ultrafeedback) show that strong LLMs excel at simulating human preferences. Consequently, numerous efforts have been made to utilize LLMs as preference data generators for scaling up.
RLAIF lee2023rlaif curates a comprehensive dataset that amalgamates the Reddit TL;DR corpus volske-etal-2017-tl, OpenAI’s Human Preferences stiennon2022learning, and the HH-RLHF bai2022training dataset, the preferences of which are annotated using PALM 2 instead of human. The results of the experiments indicate that scaling up using AI feedback significantly enhances the model’s training performance.
Open-Hermes-Preferences open_hermes_preferences is a comprehensive dataset containing roughly 1 million AI-generated preferences. It integrates outputs from this dataset and two additional models and PairRM (jiang2023llmblender) is employed as the preference model for evaluation and ordering of responses.
ULTRAFEEDBACK cui2023ultrafeedback employs GPT-4 to develop ULTRAFEEDBACK, an expansive, superior-quality, and varied preference dataset designed to overcome the scarcity and constraints of existing preference data.
UltraChat ding2023enhancing is a million-scale multi-turn instructional conversation dataset. Unlike datasets built around specific tasks, UltraChat encompasses a wide array of human-AI interaction scenarios. It leverages advanced techniques like meta-information, in-context expansion, and iterative prompting, alongside two separate ChatGPT Turbo APIs for realistic and informative conversation generation.
5 Feedbacks
Figure 6: The illustration of the reward received by the model during the preference learning. For a data sample $(x,\hat{y})$, where the $\hat{y}$ is unlabeled candidate output, the reward function is supposed to provide the feedback, which can be the reward score $r$ or the preference label. According to whether we need to train a specific reward model, the reward function can be categorized into direct reward and model-based reward.
In this section, we elaborate on the preference feedback received by the model in preference learning. Following shao2024deepseekmath, the feedback in this paper refers broadly to the preference indicators that can influence the gradient of the model during the training process. Herein, it can not only serve exactly as the reward in the methods using reinforcement learning but also be preference labels or other feedback utilized by algorithms that do not explicitly employ reinforcement learning. Formally, given a data instance $(x,\{\mathbf{\hat{y}\}})$, where $\{\mathbf{\hat{y}\}}={\hat{y}_{1},\hat{y}_{2},...,\hat{y}_{i}}$ and $i\geq 1$, the environment aligned with human preference is supposed to give out the reward, which could be preference $y_{i}>y_{j}$ or a scalar $r$. As shown in Figure 6, we survey various types of feedback in preference learning, categorizing them into two classes: direct feedback, and model-based feedback.
5.1 Direct Feedback
Direct feedback refers to the feedback that can be directly obtained without training a specific reward model.
Labeled Datasets
One of the most direct ways to obtain feedback is through labeled datasets annotated by humans. The labeled preferences within the dataset can be directly utilized for model training in offline methods. We cover the recent advancement of the existing datasets on preference learning in Section 4.2.
Hand-designed Rules
The other way to obtain direct reward is to use hand-designed rules as a reward. Due to the specificity of the rules, it is difficult to establish a unified criterion that encompasses all methods. Different tasks may adhere to different sets of rules.
For the task of mathematical reasoning, rft utilize the correctness of the reasoning paths as a metric to control the training data. Following shao2024deepseekmath, which provides another perspective of these series of methods, the reward can be calculated by $r=\mathbb{I}(c)$, where the reward equals to 1 if the COT reasoning path is correct and 0 otherwise. xin2024deepseekproveradvancingtheoremproving; xin2024deepseekproverv15harnessingproofassistant obtaining the feedback of the math theorem prover using the automated proofing tool (LEAN (LEAN_4)). For machine translation, xu2024contrastive utilize the results of the reference-free QE model to obtain the preference of different translation candidates and further optimize the model using CPO, an improvement of the DPO algorithm. For code generation, shen2023pangucoder2 rank the model output according to unit tests and heuristic preferences. For each data, they assign different scores from low to high based on the different situations of the test results. The preference obtained from the current ranking directly affects the final training loss of the model. liu2023rltf and dou2024stepcoder convert the results of unit tests under different scenarios into scalars using hand-designed rules and further optimize the model using RL algorithms. For summarization, gao2024aligning explores interactive learning for the model by using textual human edits on the agent’s outputs, which proved to be simple and effective.
5.2 Model-based Feedback
In this section, we conduct a survey of model-based feedback, encompassing reward signals from reward models, pairwise scoring models, and LLM-as-a-Judge feedback.
5.2.1 Reward Model
Training a reward model is predicated on constructing a classifier that can anticipate human preference probability, $p$, between two outputs.
Bradly-Terry based Reward Model
One line of research models the preference of humans using the Bradley-Terry modelbradley1952rank. This involves training the model to estimate $p$, derived from the comparison between two potential responses, by maximizing the likelihood of the preferred output. Parameters of this model are optimized through a loss function that emphasizes the difference in preference between a chosen and a rejected output:
$p^{*}(y_{1}\succ y_{2}\ |\ x)=\frac{\text{exp}(r^{*}(x,y_{1}))}{\text{exp}(r^{*}(x,y_{1}))+\text{exp}(r^{*}(x,y_{2}))}.$ (2)
The model is typically optimized by a negative log-likelihood loss:
$\mathcal{L}_{r}=-\log\sigma\left(r^{*}\left(y_{c},x\right)-r^{*}\left(y_{r},x\right)\right)$ (3)
where $y_{r}$ represents the rejected output and $y_{c}$ represents the chosen output. At inference time, the reward model returns a scalar $p^{*}(y_{1}\succ y_{2}\ |\ x)$ which represents the probability that the output would be the preferred response.
Binary Classifier based Reward Model
For tasks where the quality of a case can be determined by its outcomes, directly labeling samples to train a binary classifier as a reward model is a straightforward and stable approach. For instance, in mathematical reasoning, a sample can be labeled based on whether the response yields the correct final answer. Similarly, in code generation tasks, labeling can be done by checking if the generated code passes specified tests. Unlike tasks such as text summarization or dialogue generation, which require pairwise comparisons of examples, these direct assessment methods simplify the preference labeling process.
Unlike the traditional Bradley-Terry Reward Model, once the labels for the data are obtained, the reward model can be trained using point-wise binary classification loss without needing to construct pairwise data. The BCE training loss is as follows:
$L_{r}=-[rlog(\hat{r})+(1-r)log(1-\hat{r})]$ (4)
where $r$ is the preference label and $\hat{r}$ is the predicted reward.
RM Training Optimization
In pursuit of a better reward model, numerous studies optimize the existing reward models from various perspectives.
One line of research seeks to obtain better preference data. lee2023rlaif capitalizes on the capabilities of off-the-shelf LLMs to generate preference labels, potentially decreasing the need for expensive and time-consuming human annotation. The study showcases that RLAIF can reach or even surpass the performance levels of RLHF across multiple tasks. jinnai2024regularized explore using Kullback–Leibler divergence and Wasserstein Distance to regularize the Best-of-N sampling, which proved to be effective in mitigating the reward hacking problem during reward modeling. pace2024westofn utilizes West-of-N to generate better synthetic preference data, extending Best-of-N sampling strategies from language model training to the reward model training.
Another line of research focuses on model ensembling to improve the overoptimization and uncertainty estimation of the reward model. coste2024reward use reward model ensemble to mitigate the reward model overoptimization. zhai2023uncertaintypenalized consider LoRA-based ensemble, while their work focuses on an uncertainty penalized objective in RL-finetuning. ramé2024warm consider a different approach of averaging the weights of multiple reward models instead of ensembling their predictions. zhang2024improving explore multiple ensembling methods for developing efficient ensemble approaches.
Exploring another dimension, research on fine-grained rewards is gaining momentum. NEURIPS2023_b8c90b65 introduce Fine-Grained RLHF, a framework that enables training and learning from reward functions that provide rewards in multiple aspects after every segment. In contrast to outcome supervision, which provides feedback for a final result, uesato2022solving, lightman2023lets and yu2024ovm explore process supervision, which provides reward for each intermediate reasoning step. However, the training data for PRM is constrained by the high effort required for annotation, and how to efficiently construct step-level training data remains a challenge. wang2024mathshepherd construct process supervision data in an unsupervised manner, which proved to be effective for mathematical reasoning.
Additionally, optimizing the training process of reward models is a focus area, zhou2024prior; dong2024rlhf proposes using prior constraints to mitigate the uncontrolled scaling of reward scores during training the reward model. gao2024llmcriticshelpcatch proposes a two-stage training paradigm, utilizing natural language feedback to stimulate the evaluation ability of the mathematical reward model.
5.2.2 Pair-wise Scoring Model
In addition to specially trained reward models, lightweight pairwise scoring models are widely used to provide preference signals for models jiang2023llmblender. Generally speaking, pairwise scoring models employ a specialized pairwise comparison method to distinguish subtle differences between candidate outputs. Since it is easier and more consistent to discriminate among multiple candidates rather than scoring individual candidates each time, pairwise scoring models are usually smaller and achieve better results. For instance, the PairRanker jiang2023llmblender, with only 0.4B parameters, exhibits the highest correlation with ChatGPT-based ranking and is widely used in works such as SPPO sppo and SimPO meng2024simpo. However, pairwise methods cannot provide a global score, and the number of candidates they can process at one time is limited. Consequently, obtaining a global ranking among multiple candidates or a general reward signal often incurs a higher cost.
5.2.3 LLM-as-a-Judge
A more direct and easily adjustable approach is to use LLM scoring to provide rewards for preference learning or evaluation, termed LLM-as-a-Judge. For larger models, such as GPT-4, we can specify scoring rules directly in the prompts, allowing the model to score generated responses. Extending this method further, we can implement LLM self-rewarding. For example, recent self-rewarding mechanisms yuan2024self have shown that LLMs can improve by evaluating their own responses instead of relying on human labelers. However, model judgment may introduce errors or biases. To address this issue, wu2024meta introduce a novel Meta-Rewarding step, where the model assesses its own judgments and uses that feedback to refine its judgment skills. This unsupervised approach makes the scores given by the LLM more accurate. For tasks involving complex reasoning steps, LLM-as-a-Judge often underperforms the trained scoring verifiers. To mitigate this issue, mcaleese2024llmcriticshelpcatch train a critic model prompted to accept a (question, answer) pair as input and output a plain text “critique” that points out potential problems in the answer for code generation. zhang2024generativeverifiersrewardmodeling train a generative verifier to leverage the text-token prediction capabilities of LLMs for mathematical reasoning.
6 Algorithms
Preference learning algorithms update LLMs to align with human preferences based on data and rewards. In this section, we categorize the preference algorithms into four groups: point-wise methods, pair-wise contrasts, list-wise contrasts, and training-free alignment. For some representative algorithms belonging to pair-wise/list-wise contrasts, we also provide their detailed loss designs in Table 1.
6.1 Point-wise Methods
Point-wise methods optimize the model based on a single data point $(x,y)$. Due to the fact that these methods do not require paired preference data during optimization, the cost of labeling preference data is significantly reduced. The point-wise methods are easy to implement and have demonstrated effectiveness in a series of situations (rft; star; raft; schulman2017proximal; shao2024deepseekmath; li2023remax; ethayarajh2024kto).
The simplest point-wise optimization method is the rejection sampling fine-tuning. This approach first selects high-quality data points using a reward function or rules and then fine-tunes LLMs on these selected data. The object of rejection sampling fine-tuning is shown in Eq. 5, where $y^{+}$ is the response with high reward.
$L_{RS}=-\sum_{t}\log\pi_{\theta}\left(y^{+}_{t}\mid x,y^{+}_{<t}\right)$ (5)
There are several works that demonstrate the effect of rejection sampling fine-tuning. Raft (raft) employs a reward model to rank generated samples, filtering out those that best align with human preferences and values. Using these curated samples, the model undergoes fine-tuning to enhance its friendliness and accessibility to humans. Star (star) iteratively utilizes a limited selection of rationale examples alongside a substantial dataset lacking rationales, enhancing the capacity for increasingly complex reasoning without relying on a reward model. Yuan et al. (rft) employ rejection sampling fine-tuning to enhance the mathematical reasoning capabilities of LLMs, as they discover that the chosen samples encompass a greater variety of distinct reasoning paths, which proves advantageous for tackling math problems. Although simple and straightforward, rejection sampling fine-tuning fails to leverage the data with low reward, which prevents it from learning from non-preferred data and optimizing the model further.
Proximal Policy Optimization (PPO) (schulman2017proximal) from OpenAI is one of the most representative and successful point-wise optimization algorithms. Notably, the most successful applications like ChatGPT and GPT-4 are produced by the PPO method. During the PPO optimization phase, we update the LM to maximize the return from the learned reward function $r$ using the following principle:
$\max_{\theta}J_{r}(\theta)=\max_{\theta}\sum_{x}\mathbb{E}_{a\sim\pi_{\theta}(\cdot|x)}\left[r(x,y)-\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{ref}(y|x)}\right],$ (6)
where $\pi_{ref}$ is the supervised fine-tuned model and $\pi_{\theta}$ is initialized as $\pi_{ref}$. $\beta$ is the KL divergence coefficient, which controls the deviation from the original model. Despite the impressive performance demonstrated by OpenAI, the PPO algorithm requires extensive computational resources and suffers from sample inefficiency. Another drawback of the PPO algorithm is its training instability, making it challenging to determine the appropriate hyperparameters for the PPO method. Some alternative point-wise methods have introduced enhancements to the PPO framework. For instance, ReMax (li2023remax) and GRPO (shao2024deepseekmath) eliminate the use of the critic model that PPO relies on, thereby reducing the consumption of storage resources during training.
Another point-wise method is KTO from Ethayarajh et al. (ethayarajh2024kto). This approach requires very few predetermined hyperparameters, ensuring stable training while also being resource-efficient. Due to the adoption of Kahneman & Tversky’s prospect theory, KTO doesn’t need the pair-wise preference dataset. It only requires a label denoting whether the response is preferred or not. KTO directly maximizes the utility of generations instead of maximizing the likelihood of preferences. Leveraging these point-wise preference data, KTO can achieve prior or comparable performance compared with DPO. At the same time, the KTO algorithm also demonstrates good performance in cases of extreme data imbalances (ethayarajh2024kto), which makes it a good choice in certain specific circumstances.
Algorithms Formulations of Loss Function Notes
DPO (rafailov2024direct) $-\log\sigma\left(\beta\log\frac{\pi(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}-\beta\log\frac{\pi(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\right)$
$\beta$ is a hyperparameter,
typically retained by
other DPO-like methods.
IPO (azar2023general) $\left(\log\frac{\pi(y^{+}|x)\pi_{\text{ref}}(y^{-}|x)}{\pi(y^{-}|x)\pi_{\text{ref}}(y^{+}|x)}-\frac{1}{2\tau}\right)^{2}$
$\tau$ is a hyperparameter
determining the upper bound
of the scoring margin.
$f$-DPO (wang2023beyond) $-\log\sigma\left[\beta f^{\prime}\left(\frac{\pi(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}\right)-\beta f^{\prime}\left(\frac{\pi(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\right)\right]$
$f^{\prime}$ is the derivative of
the chosen $f$-divergence.
EXO (ji2024towards) $-\sum_{y\in(y_{w},y_{l})}\frac{\left(\frac{\pi(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta_{\pi}}}{\sum_{y^{\prime}\in(y_{w},y_{l})}\left(\frac{\pi(y^{\prime}|x)}{\pi_{\text{ref}}(y^{\prime}|x)}\right)^{\beta_{\pi}}}\log\frac{\left(\frac{\pi(y|x)}{\pi_{\text{ref}}(y|x)}\right)^{\beta_{\pi}}/|\mathbf{1}_{\{y=y_{w}\}}-\epsilon|}{\sum_{y^{\prime}\in(y_{w},y_{l})}\left(\frac{\pi(y^{\prime}|x)}{\pi_{\text{ref}}(y^{\prime}|x)}\right)^{\beta_{\pi}}}$
$\beta_{\pi}$ and $\epsilon$ are hyperparameters.
$\epsilon$ is used to soften the binary
human-crafted rewards.
DPO-positive (pal2024smaug) $-\log\sigma\left(\beta\log\frac{\pi(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}-\beta\log\frac{\pi(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)}\right)-\lambda\max\left(0,\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi(y^{+}|x)}\right)$ $\lambda$ is a hyperparameter.
ORPO (hong2024orpo) $\mathcal{L}_{\text{sft}}-\lambda\log\sigma\left[\log\frac{\pi(y^{+}|x)(1-\pi(y^{-}|x))}{\pi(y^{-}|x)(1-\pi(y^{+}|x))}\right]$ $\lambda$ is a hyperparameter.
SimPO (meng2024simpo) $-\log\sigma\left[\frac{\beta}{|y^{+}|}\log\pi(y^{+}|x)-\frac{\beta}{|y^{-}|}\log\pi(y^{-}|x)-\gamma\right]$
$\gamma$ is a hyperparameter to
control the margin between
the scores of $y^{+}$ and $y^{-}$.
RRHF (yuan2024rrhf) $\mathcal{L}_{\text{sft}}+\sum_{i>j}\max\left[0,\pi(y^{i}|x)-\pi(y^{j}|x)\right]$
PRO (song2023preference) $\beta\mathcal{L}_{\text{sft}}-\sum_{k=1}^{n-1}\log\frac{\exp\left[\frac{\pi(y^{k}|x)}{1/(r^{*}(x,y^{k})-r^{*}(x,y^{n}))}\right]}{\frac{\pi(y^{k}|x)}{1/(r^{*}(x,y^{k})-r^{*}(x,y^{n}))}+\sum_{i=k+1}^{n}\exp\left[\frac{\pi(y^{i}|x)}{1/(r^{*}(x,y^{k})-r^{*}(x,y^{i}))}\right]}$
$\beta$ is a hyperparameter to
balance $\mathcal{L}_{\text{sft}}$ and the rest
of list-wise contrasts.
$r^{*}$ is an external reward model.
Table 1: Demonstrations of loss function design for different algorithms. Due to the limitation of page width, there are just several algorithms selected here for they can be placed in one line. Therefore, we strongly recommend the direct reference of their papers for more works mentioned in this survey.
6.2 Pair-wise Contrasts
liu2024chain point out that point-wise methods either solely rely on the positive candidates, learning to conduct mindless intimation instead of truly understanding human preference against those negative candidates, or they face significant optimization challenges, complicating their practical application. As a result, they propose Chain-of-Hindsight (CoH) where a pair of positive and negative candidates $y^{+}$ and $y^{-}$ are both placed in the context during fine-tuning, accompanied by corresponding prompts. This helps the tuned LLM learn from semantic contrasts. For inference, the LLM is triggered by the positive prompt (e.g. A helpful response is:) to generate preferred responses.
However, such prompt methods are insufficient to force LLM to sense distinctions about human preference. Instead, more researchers choose to manipulate its inner states (e.g. probability of generating candidates) to explicitly construct pair-wise contrastive learning. For instance, zhao2023slic utilize the formulation of SLiC (zhao2022calibrating) to apply pairwise contrasts between $y^{+}$ and $y^{-}$. One of the highlights is their expansion of the training dataset to include additional candidate $y$ sampled from the initial SFT checkpoint. Another representative method is Direct Preference Optimization (DPO) designed by rafailov2024direct. They change the objective of RLHF as an equation between the given reward model $r$, reference model $\pi_{\text{ref}}$, and corresponding optimal policy $\pi_{r}$,
$r(x,y)=\beta\frac{\pi_{r}(y\mid x)}{\pi_{\text{ref}}(y\mid x)}+\beta\log Z(x)$ (7)
where $x$ and $y$ are the context and its candidate response, respectively. Combining Equation 7 with the Bradley-Terry bradley1952rank loss used in reward models, DPO formulates a new objective for direct optimization of policy $\pi$ while containing an implicit reward learning process,
$\mathcal{L}_{\text{DPO}}(\pi;\pi_{\text{ref}})=$ $-\log\sigma(\beta\log\frac{\pi(y^{+}\mid x)}{\pi_{\text{ref}}(y^{+}\mid x)}-\beta\log\frac{\pi(y^{-}\mid x)}{\pi_{\text{ref}}(y^{-}\mid x)})$ (8)
Despite its effectiveness, azar2023general demonstrates that DPO can easily overfit the pair-wise annotations from the provided preference datasets. It arises from that DPO transforms the score of $y^{+}$ to an unbounded range using a non-linear mapping $\psi(q)=\log\frac{q}{1-q}$. This leads to an extreme optimization of the score difference while reducing the impact of the regularization term in RLHF simultaneously. The authors accordingly propose Identity-PO (IPO) that replaces the unbounded mapping with the identity mapping. In essence, IPO constrains the upper bound of the above score difference to alleviate overfitting.
After the appearance of IPO, more researchers attempt to modify $\mathcal{L}_{\text{DPO}}$ for better performance. wang2023beyond point out the constraint of reverse KL divergence in RLHF for the diversity of generated content, which can be mitigated by other $f$-divergences. They consequently abstract a general DPO-like loss function, providing a plug-and-play form for different $f$-divergences. ji2024towards claim that, unlike RLHF, DPO essentially optimizes a forward-KL divergence $\text{KL}(\pi_{\text{ref}}||\pi)$. As a substitute, they directly build the objective equivalent to the reverse-KL divergence by a simple estimation of the partition function. chen2024mallows and ramesh2024group both focus more on the input context. chen2024mallows rely on the condition of Mallow (mallows1957non) formulation to model the effect of input context on the finally acquired reward, which replaces the original reward modeling in DPO, while ramesh2024group utilize the context to place fine-grained controlling information. Moreover, some researchers find that DPO tends to reduce the log-likelihood of $y^{+}$, which could encourage LLM to generate sub-optimal responses (feng2024towards; yan20243d). pal2024smaug then proposed the DPO-positive method that appends a penalty term to mitigate this phenomenon. yu2024direct also takes the way of appending penalty term, which, however, sources from prompting the LLM itself.
Another direction is the modification of the training pipeline. Typically, DPO involves two progressive stages: the first is SFT, and the next is contrastive alignment. Hence, one way of the modification observed is to reduce the cost of fine-tuning. hong2024orpo append a term of Odds Ratio to the initial SFT loss for enhanced supervision. This design substantially follows the spirit of pair-wise contrasts, while combining the above two stages into one to shorten the pipeline. Besides, a concise DPO-like algorithm has recently been proposed, named SimPO(meng2024simpo). It inherits the framework of DPO, but eliminates the reference model $\pi_{\text{ref}}$ in the second stage to directly optimize the log-likelihood of $y^{+}$ exceeding other candidates, which, on the other hand, aligns with the nature of maximized log-likelihood of sequences in LLM inference. Other attempts can be transferring the offline DPO into online settings. For example, kim2024sdpo and gorbatovski2024learn share a motivation of dynamically updating $\pi_{\text{ref}}$ in DPO, including full replacement and soft merging. kim2024sdpo provide convincing evidence of this manner: through transforming its loss, DPO can be viewed as optimizing $\pi$ to keep $\frac{\pi(y^{+}|x)}{\pi(y^{-}|x)}$ away from $\frac{\pi_{\text{ref}}(y^{+}|x)}{\pi_{\text{ref}}(y^{-}|x)}$, and updating $\pi_{\text{ref}}$ iteratively force $\pi$ to converge to the optimal policy. morimura2024filtered share the idea of online data collection and selection with raft, but leverage it on DPO. This process aims to alleviate the effect of original low-quality data. liu2023statistical and zhang2024ts complete similar targets with more complex frameworks.
Some works also promote the application of preference alignment. For instance, hejna2023contrastive derives an offline policy LLM learning method based on the regret-based model of human preference, named Contrastive Preference Learning, which can model the preference in more complex scenarios, like robotics. With careful re-definitions of preference in specific domains, she2024mapo and lyu2024knowtuning successfully transfer DPO to multilingual reasoning and knowledge-aware QA, while zhou2023beyond, guo2024controllable and badrinath2024hybrid focus on the adaptation of multi-objective preference alignment.
6.3 List-wise Contrasts
Extending pair-wise contrasts to list-wise ones is also a natural inspiration, whose effectiveness has been demonstrated by song-etal-2024-scaling. RRHF (yuan2024rrhf) pioneers the early adoption of expanding two candidates ${y^{+},y^{-}}$ to a longer list ${y_{i}}$ using external LLM. Nevertheless, this method pairs each candidate $\{y^{i}\}$ with its inferior ones $y^{>i}$ to create multiple pairs, hence maintaining the utilization of pair-wise contrasts. song2023preference further propose Preference Ranking Optimization (PRO) to recursively applies multiple list-wise contrasts between $y^{i}$ and $y^{>i}$. hong2023cyclealign then leverage list-wise contrasts on the distillation of superior black-box LLMs, acquiring improved performance.
Vanilla list-wise contrasts may lead to performance degradation due to biased estimation of candidates, requiring more fine-grained design. wang2023making and mao2024don implement multiple calibration strategies on computed scores for different ranking objectives to alleviate over-fitting. Differently, liu2024lipo and zhu2024lire opt to design re-weighting mechanisms on each term in loss functions with external scoring information, which are beneficial to precise scoring ability of $\pi$.
6.4 Training-Free Alignment
Training-free alignment refers to approaches that do not fine-tune the language model itself, by which parameters of the language model remain unchanged after alignment. Instead, training-free alignment enhances the model output to better align with preferences by optimizing input prompts (§6.4.1) or optimizing at the output stage (§6.4.2). Optimization on the input side includes incorporating in-context learning examples (song2024icdpo; lin2023unlocking), retrieval-augmented content (xu2023align) into the prompts, and employing a prompt rewriter (cheng2023black) to refine the prompts before fed into the model. On the output side, optimization involves redistributing the model’s output probability distribution over vocabulary (yang2021fudge; deng2023reward; huang2024deal), backtracking and regenerating when harmful content is encountered during decoding (li2023rain), and adding a module after the model to rewrite the originally generated content (ji2024aligner).
6.4.1 Input Optimization
lin2023unlocking analyzes the token distribution shifts between the content generated by base models and aligned models, finding that most shifts occurred with stylistic tokens. Building on this insight, they employ system prompts and restyle the responses of in-context learning examples to align the model. xu2023align aligns the generated contents by retrieving norms most relevant to the given prompt. BPO (cheng2023black) posits that an effective prompt can lead to better responses. Based on this concept, BPO trains a sequence-to-sequence model to act as a prompt rewriter by using low-quality and high-quality prompt pairs generated by ChatGPT and uses it to optimize the input during inference.
6.4.2 Output Optimization
Paraphrasing
Aligner (ji2024aligner) trains an additional alignment module. During inference, the instruction is first fed into the original model to generate an unaligned response; this unaligned response, along with the instruction, is then fed into the alignment module to produce an aligned response.
Logits Manipulation
FUDGE (yang2021fudge), mudgal2023controlled; deng2023reward; liu2024decoding achieve alignment by modifying the model’s output probability distribution during the decoding phase, and they consequently increase the likelihood of generating aligned responses and decrease the probability of harmful responses.
Searching
RAIN (li2023rain) achieves alignment by rewindable decoding, where harmful tokens are discarded while helpful tokens are preserved, and the quality of the token is evaluated by the LLM itself. huang2024deal replace the self-evaluation with customized reward models to allow more fine-grained requirements of personal preference in aligned decoding. ICDPO (song2024icdpo) designs a two-stage Best-of-N-like process to optimize output, where the final response is selected among multiple candidates sampled from a local LLM upon In-context Learning (ICL), according to their different degrees of human preference. Unlike the traditional Best-of-N that relies on external verifiers (e.g. reward models) for response selection, ICDPO proposes a skillful formulation that aggregates the states before and after ICL as a joint estimation, which is completed solely by the local LLM itself. It can achieve comparable performance as fine-tuning but requires just a few high-quality demonstrations, reducing the implementation cost.
7 Evaluation
The methods for evaluating alignment are divided into two parts: rule-based evaluation as Sec. 7.1 and LLM-based evaluation as Sec. 7.2.
7.1 Rule-based Evaluation
Rule-based evaluation is generally conducted in the scheme where the dataset has ground-truth output for each input. In this way, the evaluation can be done by using the widely used automatic metrics including Accuracy, F1, Exact-Match rajpurkar2016squad and ROUGE lin2004rouge.
Current evaluation of general-purpose LLM mainly focuses on evaluating some core tasks before they can be generalized to satisfy various practical needs. LLMs achiam2023gpt; touvron2023llama; touvron2023llama2; vicuna2023 are evaluated on a multi-aspect evaluation set to cove key capabilities.
Factual Knowledge
Factual Knowledge is essential for language models to serve users’ information needs, including Massive Multitask Language Understanding dataset (MMLU) hendrycks2020measuring, C-Eval huang2024c and Massive Multitask Language Understanding in Chinese (CMMLU) li2023cmmlu.
Math
Mathematical reasoning includes the test split of Grade School Math dataset (GSM8K) cobbe2021training, MATH hendrycks2020measuring and Chinese Elementary School Math Word Problems dataset (CMATH) wei2023cmath.
Reasoning
Reasoning is a fundamental ability for LLMs, especially to solve complex problems, including Big-Bench-Hard (BBH) suzgun2022challenging.
Closed-Book Question Answering
Closed-Book question answering includes TriviaQA joshi2017triviaqa, NaturalQuestions kwiatkowski2019natural, CSQA saha2018complex and StrategyQA geva2021did.
Coding
Coding is a special application that people tend to utilize LLMs and might be significant for integrating LLMs with external tools. LLMs would be better at tool usage and function calling with better coding skills. Coding benchmarks includes MBPP austin2021program and HumanEval chen2021evaluating. Nowadays Coding evaluation benchmarks tend to evaluate LLMs at repo-level code generation including SWE-Bench jimenez2023swe and ML-Bench liu2023ml.
However, the rule-based evaluation strategy with standard metrics suffers from significant drawbacks. The standard metrics only present whether the models’ outputs are close to the ground truth outputs, but the currently popular tasks are generally open-ended (i.e., summarization). Calculating standard metrics between models’ outputs with the ground truth labels is a misleading evaluation.
7.2 LLM-based Evaluation
With the recent advances of LLMs, LLM-based assistants have started to exhibit artificial general intelligence across diverse tasks (i.e., writing, chatting, and coding). Rule-based evaluation of the aforementioned tasks is challenging. Some recent works gudibande2023false; chiang2024chatbot have found inconsistencies between the performance of language models in human evaluations and NLP benchmark, which may be due to existing evaluation (rule-based evaluation) that only measures LLMs’ core capability on a confined set of tasks (e.g., multi-choice knowledge or retrieval questions) without considering the alignment with human preference in open-ended tasks. There is an emergent need for a robust and scalable automated method to evaluate LLM alignment with human preferences. Consequently, Using Large Language Models (LLMs) as proxies for human evaluation to assess the quality of other LLMs has emerged as a cost-effective and promising approach.
7.2.1 LLM-based Evaluation Methods
LLM-based evaluation methods can be mainly categorized into the following three types:
Pairwise Comparison
LLM evaluators are provided with instructions and the corresponding outputs from two models, then asked to choose the preferred one or declare a tie. wang2023large; wang2023automated; li2023generative; wang2023pandalm; kim2023prometheus Pairwise Comparison is generally the most prevalent method and boasts the highest consistency between human evaluation and LLM-based evaluation, but this method itself is not scalable because the computational costs (.e., the number of possible pairs) will grow significantly as the number of evaluated models increases. AlpacaEval alpaca_eval is a fast, cost-effective, and reliable LLM-based automatic evaluation system. It uses the AlpacaFarm dubois2023alpacafarm evaluation set, which is designed to assess models’ ability to follow general user instructions. Model responses are compared to reference responses using GPT-4-based auto-annotators, producing the win rates presented above. AlpacaEval alpaca_eval exhibits a high level of agreement with human annotations, and its leaderboard rankings strongly correlate with those generated by human evaluators. AlpacaEval introduces a metric based on the win rate of two LLM-generated responses, as judged by human evaluators. AlpacaEval 2.0 dubois2024length further refines this by introducing a length-controlled win rate, which debiases the win rates by accounting for output length differences.
Single Answer Grading
Alternatively, an LLM evaluator can be asked to assign an evaluation score to a single instruction and the corresponding answer. liu2303g; jain2023multi; kocmi2023large; li2023prd; wang2023automated; kim2023prometheus Single Answer Grading is efficient for ranking multiple models but fails to discern subtle differences between specific pairs and may show significant score fluctuations across evaluations wang2023large; li2023prd; zheng2024judging; zeng2023evaluating.
Reference-Guided Grading
Providing reference answers is crucial for evaluating models in tasks with objective human preferences, such as mathematics, translation, etc. wang2023large; wang2023pandalm; kim2023prometheus However, this method requires high-quality annotations for reference answers.
While this method is currently the most prevalent and boasts the highest human consistency, it suffers from scalability issues as the number of models to evaluate increases, resulting in quadratic growth in possible pairs and consequently escalating computational costs  zeng2023evaluating.
7.2.2 LLM-based Evaluation Models
The most-preferred model as an LLM-based evaluator is the priority models including GPT-4, which may be motivated by the fact that these models are typically trained using RLHF to align with human preferences and have demonstrated strong human consistency. bai2022training Intuitively, evaluators should be capable of distinguishing between good and bad responses  liu2303g; liu2303g; kocmi2023large; chiang2023can; wang2023large; li2023prd. Utilizing state-of-the-art models in this way leads to outstanding performance and broad generalizability. However, drawbacks include high costs and the potential irreproducibilityİn response, recent research has shifted towards fine-tuning smaller, open-source LLMs for evaluation purposes, aiming to achieve performance close to GPT-4. These models are primarily created by meticulously constructing high-quality evaluation data and fine-tuning open-source models. Compared with employing priority models, Fine-tuning small models enhances the model’s evaluation capability in certain aspects, mitigating the potential of bias (i.e. position bias) and significantly reducing costs. However, experiment results indicate that while the fine-tuned evaluation models achieve superior accuracy on their respective in-domain test sets, they still exhibit limitations, including a lack of generalization, overfitting on specific evaluation strategies, and a bias towards surface quality.
7.2.3 Limitations
The LLM-based evaluator has been found to exhibit certain biases and limitations.  wang2023large note that LLM-based evaluator inevitably has position bias (i.e., when using GPT-4 for pairwise comparison, the evaluation results can be easily hacked by altering the order in which candidate answers appear in context). Another bias of LLM-based evaluator is that the LLM-based evaluator exhibits a tendency to prefer more verbose outputs zheng2024judging, shows a predisposition towards outputs generated by models similar to itself  gudibande2023false; zheng2024judging, and displays a limited capability in evaluating subjects like mathematics, reasoning, and other areas that still pose challenges for LLMs zeng2023evaluating. To systematically quantify the performance of LLM-based Evaluators, several works have introduced meta-evaluation benchmarks. FairEval wang2023large, MT-Bench zheng2024judging, and LLMEval zhang2023wider assess whether LLM-based evaluators demonstrate high agreement with humans by utilizing manually annotated preference datasets.  zeng2023evaluating proposed a meta-evaluation benchmark called LLMBar, which includes an Adversarial set. Notably, all models, including GPT-4, struggled on the adversarial set without the use of additional strategies.
8 Future Directions
Better quality and more diverse preference data.
In the preference learning scenario, the final performance of the model to a large extent depends on the quality and diversity of the preference data (guo2024direct; song-etal-2024-scaling). Therefore, further research can be conducted on this domain. For example, synthetic data techniques can be utilized to ensure prompt quality (yuan2024self; pace2024westofn). Besides, advanced sampling techniques may be explored to enhance the sampling diversity and quality of the model response (xie2024montecarlotreesearch).
Reliable feedback and scalable oversight.
The optimization objective of preference learning comes from the feedback, and thus reliable feedback plays an important role. Some reliable feedback such as code compiler (shen2023pangucoder2; grubisic2024compilergeneratedfeedbacklarge) or proof assistant (xin2024deepseekproverv15harnessingproofassistant) is explored, but they are limited to code or math domain. It would be valuable if we could extend them into more general domains. In addition, more research is required in cases where humans cannot provide reliable feedback anymore to enable scalable oversight for the next-generation super-intelligence, such as recursive reward modeling (yuan2024self), or weak-to-strong technique (burns2023weaktostronggeneralizationelicitingstrong; chen2024self).
Advanced algorithm for preference learning.
Data and feedback determine the upper bound of the model performance, and a good training algorithm can help us approach this upper bound as much as possible. In the future, better training algorithms should strive to meet the following requirements: (1) better approach the performance upper bound; (2) more robust to the provided data and feedback (liang2024roporobustpreferenceoptimization; ramesh2024group); (3) higher training efficiency and therefore can be scaled up (li2023remax; meng2024simpo). In fact, there are already many optimized variants of PPO and DPO for preference learning. However, the performance of these algorithms may be inconsistent across different models and task settings (saeidi2024insightsalignmentevaluatingdpo). Finding the most effective variant from a theoretical perspective is also a very practical topic, which we leave to our future work.
More comprehensive evaluation for LLM.
The existing evaluation datasets are not comprehensive enough to assess the capabilities of models, and the form of the questions is also relatively homogeneous (e.g., multiple-choice questions). Although more and more open-ended generation evaluation benchmarks are proposed, factors such as evaluation bias (wang2023large) and the cost of evaluation (biderman2023pythiasuiteanalyzinglarge) still trouble us. We need more comprehensive, reliable, and diverse evaluation methods and benchmarks, which are complementary to the development and progress of large language models.
9 Conclusion
In this survey, we decompose the strategies of preference learning into several modules: Model, Data, Feedback, and Algorithm. By distinguishing different strategies according to their variants, we build a unified view of preference learning strategies and establish connections among them. We believe that although the core objectives of these alignment algorithms are essentially similar, their performance can vary significantly across different application scenarios. We leave the exploration of which variants perform better in specific contexts as our future work. Finally, we hope this survey provides researchers with a further understanding of preference learning and thereby inspires further research in this field.
[…]
[heading=bibintoc]
Appendix A The discussion about Online vs. Offline
Unlike traditional reinforcement learning, discussions about Online and Offline in preference learning of large language models lack a unified definition. Essentially, the distinction between online and offline lies in whether feedback is obtained in real-time from the environment, that is, whether the policy model interacts with the environment in real-time. For preference data $(x,y,r)$, if the feedback $r$ is obtained in real-time from the environment, it is considered online. For instance, in RLHF, the feedback of on-policy data is obtained in real time from the reward model during the model training.
However, some studies (alignment_guidebook) delve deeper into whether a fixed reward model can genuinely represent the environment. These studies argue that a static reward model cannot consistently reflect the environment, thereby classifying RLHF with a fixed reward model as offline. In this review, for the sake of simplicity and clarity, we do not dwell on this distinction too much; instead, we define the output of the reward model as a form of signal from the environment. Consequently, in this context, methods like RLHF are considered online and DPO is offline.